40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 31     Entry time: Thu Jul 7 17:44:11 2016
Author: Gabriele 
Type: Facility 
Category: Daily Progress 
Subject: Vacuum chamber tests 

[Alena, Gabriele]

This afternoon we opened the tall belljar vacuum chamber, and took everything out of it. All the stuff that was inside the chamber is "temporarily" stored on the floor beside the optical table.

We installed a "spacer" into the chamber, made from one of the optical table legs that were sitting in the hallway. We installed one of the aluminum base plates on top of it, so that the optical components will be at the level of the viewport. Another leg and a thinner base plate are installed out of the chamber, at a similar level.

After this we closed the chamber with one of the flats used for the old chamber, and a rubber o-ring. We started the roughing pump, quickly reached a pressure below 1 mTorr and switched on the turbo pump. Unfortunately, it seems that the low pressure gauge is not working properly, so we don't know what's the pressure right now. We'll check the gauge and controller tomorrow morning and swap it out if needed.

ELOG V3.1.3-